# AssemAI: Interpretable Image-Based Anomaly Detection for Manufacturing Pipelines

### v1

@@ IV-A Task Description
    Cropping each image $I_{j,s}$ using its bounding box $B_{j,s}$ which is mapped to its specific state. This effectively isolates the regions of interest while removing background noise.
@@
    Applying AssemAI to classify the cropped images.
The primary objective is to detect anomalies in the images, leveraging the bounding boxes to enhance the model’s ability to focus on the critical parts of the image. To address class imbalance, we use a weighted cross-entropy loss, where higher weights are assigned to classes with fewer images. Mathematically, we aim to minimize the weighted classification error defined as:
$\min_{\theta}\frac{1}{n}\sum_{j=1}^{n}\sum_{s\in\{4,9\}}w_{y_{j,s}}\mathcal{L}(f(I_{j,s};\theta),y_{j,s})$
where $\theta$ represents the model parameters, $\mathcal{L}$ is the cross-entropy loss function and $w_{y_{j,s}}$ is the weight assigned to the class $y_{j,s}$.
By focusing on specific cycle states and using object detection to preprocess the images, we aim to improve the accuracy and reliability of our anomaly detection system. The ultimate goal is to create a robust model that accurately detects anomalies in the toy rocket manufacturing process.
@@ V Experiments
V-A Common Hyperparameters and Training Setup
@@
The preprocessed dataset was split into training (80%) and testing (20%) sets. For all models, we used Cross-Entropy Loss and the Adam optimizer [49], tuning hyperparameters like epochs, batch size and learning rate. The best model was saved based on validation accuracy. The dataset preprocessing involved resizing images to the required input dimensions and normalizing them according to the specific requirements of each model.
V-B Baselines
V-B1 Custom CNN
@@
The Simple CNN model architecture included two convolutional layers (32 and 64 filters respectively), each followed by a max-pooling layer. After flattening, the output was passed through a fully connected layer with 512 neurons, then to the final layer corresponding to the number of classes (5).
V-B2 Custom ViT
@@
The Vision Transformer (ViT) model [50] was implemented from scratch, starting with a patch embedding layer that splits the input image into non-overlapping patches, each embedded into a high-dimensional space. This was followed by a series of transformer blocks for learning spatial relationships within the image patches. Each transformer block consisted of a multi-head self-attention mechanism and a feed-forward neural network, with layer normalization and residual connections applied at each step. The final classification head mapped the output to the desired number of classes (5).
V-B3 Pretrained ViT
@@
We utilized the pre-trained Vision Transformer (ViT) model, specifically google/vit-base-patch16-224. Images were pre-processed by dividing each image into a sequence of fixed-size non-overlapping patches, then linearly embedded. A [CLS] token was added to represent the entire image, facilitating classification. Absolute position embeddings were incorporated and the resulting sequence of vectors was fed to the standard Transformer encoder. The ViTImageProcessor resized and normalized images to the required 224x224 resolution. The model was loaded with half-precision (torch.float16). Training utilized PyTorch Lightning’s Trainer, with a weighted Cross-Entropy loss function and the AdamW optimizer with a learning rate scheduler. The best checkpoint was saved by monitoring the validation loss.
V-C Proposed Approach
@@
We implemented an EfficientNet-B0 model [16] for image classification tasks. The EfficientNet-B0 architecture was selected for its state-of-the-art performance and efficiency in image classification. It leverages a compound scaling method to optimize model depth, width and resolution, achieving high accuracy with fewer parameters compared to traditional networks. The model was pre-trained on ImageNet and adapted for our task by modifying the final classification layer to match the number of classes in our dataset (5 classes). The model was trained for 10 epochs, with the best model saved based on validation accuracy.
@@ VI Results
Table IV summarizes the results of our experiments on the test set and the ablation studies across different baselines. We evaluate the performance using four metrics: weighted averages of precision, recall, F1-score and accuracy. The weighted averages are calculated based on the four types of anomaly classes and the normal class. The EfficientNet model achieves 96% overall accuracy and 95% of precision, 96% of recall, 95% of F1 score, respectively. Figure 4 depicts the performance of detecting various anomaly types and the normal class across various modeling approaches. It can be observed that among all the models, EfficientNet with the YOLO-FF module gives the best results in detecting all the types of classes.
TABLE IV: Experimental Results
OD Model
WP
WR
WF1
Accuracy Support*
Simple CNN
+YOLO-FF
92.00±
@@
*Support-Number of samples used. WP-Weighted Average Precision, WR-Weighted Average Recall, WF1-Weighted Average F1-Score
Figure 4: Experimental results. Class 1-5 denotes anomaly types: [No Anomaly], [NoNose], [NoNose,NoBody2], [NoNose,NoBody2,NoBody1] and [NoBody1] respectively
@@ VI-A1 User level explainability
In this work, we employ process ontology designed and developed for Future Factories Rocket assembly line. In comparison with conventional ontologies, process ontology not only captures the definition of sensors but also captures the procedural nature of the assembly process. This aids in understanding the involvement of sensors and equipments at a given point in assembly process. The Future Factories assembly is divided into 21 cycle states which forms the basis of process ontology construction. The specific features of process ontology are as follows: (i) consists of definition and item specification of sensors and equipments, (ii) relationship between the sensors and equipments, (iii) function and involvement of each sensors and robots with respect to the cycle states (iv) expected (or anomalous) values of sensor variables in with respect to each cycle state (v) type of anomaly that could be associated with each cycle state (iv) sensor values and other knowledge can be dynamically updated as per the change in experiment set up. Capturing the procedural nature of the assembly process aids in understanding the contribution of sensors in anomalies. For example, if the initial stage of assembly line is being analyzed for anomaly, it can be understood from ontology that Robot-4 and its corresponding sensors do not contribute to this anomaly.
The goal of process ontology is to explain and assess the output from the models. Given an image, if the model detects an anomaly, the expected values of the sensors present in that image can be obtained and provided to the user. On the other hand, process ontology can also verify if the model predictions are correct in certain cases (Figure 5). Since each image is associated with a timestamp which in turn can be mapped to cycle state, the predicted output from the model can be verified using anomaly types defined in the process ontology. Illustrating using an output from the model using Figure 5, we notice that the input image is associated with cycle state 4. The model predicted the image to be anomalous with type NoNose. This is an incorrect prediction as per the ontology as NoNose anomaly can happen only from cycle state 8 onwards. Using this knowledge and verifying the outputs model predictions, it is found that the model incorrectly predicted NoBody2 anomaly 105 out of 801 times and NoNose anomaly 245 out of 1145 times.
@@ VII Deployment of AssemAI
After training and testing the model, we then worked on deploying AssemAI on the Future Factories lab. This deployment strategy can be seen in 6. AssemAI requires two separate types of inputs to operate in real-time. The script that utilizes the model should be able to obtain the current cycle state of the assembly process. This data tag allows the script to ensure that the images captured and input into the model are from cycle states 4 and 9. This tag is obtained by connecting to the OPC UA server running on the Programmable Logic Controller (PLC) and constantly reading the tag as it is updated. The cameras used in this deployment are separate from the PLC’s network. Images taken by the cameras can be acquired by connecting directly to them through a USB 3.0 wired connection. These cameras are industry-grade Basler cameras with a Python library, pyplon, which simplifies the image-capturing process. Since this is a wired connection, we can ensure that there is minimal communication lag time between the image request and capturing process. As such, once the cycle state tag is read as either four or nine, the images are captured from the cameras and sent into the model for detection.
Figure 6: Deployment Architecture
@@ VIII Conclusion and Further Work
In this study, we curated an industry-standard dataset tailored for assembly processes using the novel YOLO-FF model and introduced AssemAI, a standardized image-based anomaly detection pipeline. We began with a simple CNN Model, which provided a baseline for anomaly detection through a straightforward architecture and standard training techniques. Building on this, we explored ViT and EfficientNet models to enhance classification accuracy and efficiency. A significant aspect of our approach was the focus on interpretability, where we aimed to understand both model and user lever explainability to high-level phenomena such as structural integrity. This not only improved model understanding but also provided actionable insights for anomaly detection. Our findings indicate that EfficientNet offers significant improvements over traditional methods. Future work should explore hybrid architectures, which take other modalities like time series and textual, further interpretability techniques and deployment on edge devices for real-time anomaly detection applications to extend these findings to other domains and production environments. Also, to improve how domain experts understand our approach, we suggest creating abstract representations of causal factors associated with anomalies. This approach focuses on linking sensor data to broader concepts like structural issues or gripper malfunctions, providing a more comprehensive view of anomalous events.
@@ Acknowledgment
This work was supported in part by NSF grants #2119654, “RII Track 2 FEC: Enabling Factory to Factory (F2F) Networking for Future Manufacturing” and SCRA grant “Enabling Factory to Factory (F2F) Networking for Future Manufacturing across South Carolina”.

### v2

@@ IV-A Task Description
    Cropping each image $I_{j,s}$ using its bounding box $B_{j,s}$, which is mapped to its specific state. This effectively isolates the regions of interest while removing background noise.
@@
    Detecting objects in the cropped images using an object detection model. This will generate the region of interest (i.e., rocket parts) and their bounding boxes.
  4. 4.
    Classifying anomalies from the detected objects using an anomaly detection model.
The main goal is to detect anomalies in the images by using bounding boxes to help the model focus on the important parts. To address class imbalance, we use a weighted cross-entropy loss, where higher weights are assigned to classes with fewer images. We aim to minimize the weighted classification error defined in equation 2:
$\min_{\theta}\frac{1}{n}\sum_{j=1}^{n}\sum_{s\in\{4,9\}}w_{y_{j,s}}\mathcal{CE}(f(I_{j,s};\theta),y_{j,s})$ (2)
where $\theta$ represents the model parameters, $\mathcal{CE}$ is the cross-entropy loss function, and $w_{y_{j,s}}$ is the weight assigned to the class $y_{j,s}$. By focusing on specific cycle states and using object detection to preprocess the images, we aim to improve the accuracy and reliability of our anomaly detection system.
@@ V Experiments
V-A Experiments with YOLO-FF Model
The YOLOv9 model is fine-tuned on the derived labeled image dataset resulting from the OWL-ViT model. It is trained for 50 epochs. During training, the model achieved an accuracy of approximately 99% on the validation dataset.
V-B Common Hyperparameters and Training Setup
@@
The derived image dataset is split into training (80%) and testing (20%) sets. We used Cross-Entropy Loss and the Adam optimizer [48] for baselines and the proposed models, tuning hyperparameters like epochs, batch size and learning rate. The best model is saved based on validation accuracy. The dataset preprocessing involved resizing images to the required input dimensions (224×224 pixels), normalizing ([0.485,0.456,0.406]) and standardizing ([0.229,0.224,0.225]) according to the specific requirements of each model.
V-C Baselines
V-C1 Custom CNN
@@
The Simple CNN model architecture includes two convolutional layers (32 and 64 filters respectively), each followed by a max-pooling layer. After flattening, the output is passed through a fully connected layer with 512 neurons, then to the final layer corresponding to the number of classes (5).
V-C2 Custom ViT
@@
The Vision Transformer (ViT) model [49] is implemented from scratch, starting with a patch embedding layer of size 16 that splits the input image into non-overlapping patches, each embedded into a high-dimensional space of size 768. This is followed by a series of transformer blocks for learning spatial relationships within the image patches. Each transformer block consists of a multi-head self-attention mechanism and a feed-forward neural network, with layer normalization and residual connections applied at each step. The final classification head maps the output to the desired number of classes (5).
V-C3 Pre-trained ViT
@@
We utilize the pre-trained Vision Transformer (ViT) model, specifically google/vit-base-patch16-224. Images are preprocessed by dividing each image into a sequence of fixed-size non-overlapping patches, then linearly embedded. A [CLS] token is added to represent the entire image, facilitating classification. Absolute position embeddings are incorporated and the resulting sequence of vectors is fed to the standard Transformer encoder. The ViTImageProcessor resizes and normalizes images to the required 224x224 resolution. The model is loaded with half-precision (torch.float16). The training utilized PyTorch Lightning’s Trainer, with a weighted Cross-Entropy loss function, and the AdamW optimizer with a learning rate scheduler. The best checkpoint is saved by monitoring the validation loss.
Figure 4: Experimental results with OWL-ViT. Class 1-5 denotes anomaly types: [No Anomaly], [NoNose], [NoNose,NoBody2], [NoNose,NoBody2,NoBody1] and [NoBody1] respectively
V-D Proposed Modeling Approach
@@
We implement an EfficientNet-B0 model [16] for image classification tasks. The EfficientNet-B0 architecture is selected for its state-of-the-art performance and efficiency in image classification. It leverages a compound scaling method to optimize model depth, width and resolution, achieving high accuracy with fewer parameters compared to traditional networks. The model is pre-trained on ImageNet and adapted for our task by modifying the final classification layer to match the number of classes in our dataset (5 classes).
@@ VI Results
Table IV summarizes the results of our experiments on the test set and the ablation studies across different baselines. We evaluate the performance using four metrics: weighted averages of precision, recall, F1-score and accuracy. The weighted averages are calculated based on the four types of anomaly classes and the normal class. The EfficientNet model achieves 96% overall accuracy and 95% of precision, 96% of recall, 95% of F1 score, respectively. Figure 4 depicts the performance of detecting various anomaly types and the normal class across various modeling approaches with the YOLO-FF object detection approach. It can be observed that among all the models, EfficientNet gives the best results in detecting all the types of classes.
TABLE IV: Experimental Results of AssemAI
Model WP WR WF1 Accuracy Support*
Simple CNN*
82.00±
@@
Models marked in * are experimented with the derived image dataset. EfficientNet model is experimented for original images (includes all cycle states) and for filtered images (includes cycle states four and nine). Bold indicates the best performance.
@@ VI-A1 User level explainability
In this work, we employ process ontology designed and developed for the Future Factories Rocket assembly line. In contrast to conventional ontologies, process ontology not only captures the definition of sensors but also captures the procedural nature of the assembly process. This aids in understanding the involvement of sensors and types of equipment at a given point in the assembly process. The Future Factories assembly is divided into 21 cycle states, which form the basis of the ontology construction process. The specific features of process ontology are as follows: (i) consists of definition and item specification of sensors and types of equipment, (ii) relationship between the sensors and types of equipment, (iii) function and involvement of each sensor and robot with respect to the cycle states (iv) expected (or anomalous) values of sensor variables in with respect to each cycle state (v) type of anomaly that could be associated with each cycle state (iv) sensor values and other knowledge can be dynamically updated as per the change in experiment set up. Capturing the procedural nature of the assembly process aids in understanding the contribution of sensors in anomalies. For example, if the initial stage of the assembly line is being analyzed for anomaly, it can be understood from the ontology that Robot-4 and its corresponding sensors do not contribute to this anomaly.
The goal of process ontology is to explain and assess the output from the models. Given an image, if the model detects an anomaly, the expected values of the sensors present in that image can be obtained and provided to the user. On the other hand, process ontology can also verify if the model predictions are correct in certain cases (Figure 5). Since each image is associated with a timestamp which in turn can be mapped to the cycle state, the predicted output from the model can be verified using anomaly types defined in the process ontology. Figure 5 illustrates that the input image is associated with cycle state 4. The model predicted the image to be anomalous with type NoNose. This is an incorrect prediction as per the ontology as NoNose anomaly can happen only from cycle stage 8 onwards. Using this knowledge and verifying the outputs of model predictions, it is found that the model incorrectly predicted NoBody2 anomaly 51 out of 801 times and NoNose anomaly 106 out of 1145 times. The ontology also acts as an additional layer that catches the misclassification of the model, enhancing the robustness of the proposed anomaly detection pipeline.
@@
Figure 6: Deployment Setup of AssemAI
Figure 7: Web User interface of AssemAI
@@ VII Deployment of AssemAI
After training and testing the model, we deploy AssemAI on the Future Factories lab. This deployment setup can be seen in Figure 6. AssemAI requires two separate types of inputs to operate in real time. The script that utilizes the model should be able to obtain the current cycle state of the assembly process. This data tag allows the script to ensure that the images captured and input into the model are from cycle states four and nine. This tag is obtained by connecting to the OPC UA server running on the Programmable Logic Controller (PLC) and constantly reading the tag as it is updated. The cameras used in this deployment are separate from the PLC’s network. Images taken by the cameras can be acquired by connecting directly to them through a USB 3.0 wired connection. These cameras are industry-grade Basler cameras with a Python library, pyplon, which simplifies the image-capturing process. Since this is a wired connection, we can ensure that there is minimal communication lag time between the image request and the capturing process. As such, once the cycle state tag is read as either four or nine, the images are captured from the cameras and sent into the model for detection. Figure 7 illustrates the web user interface of the deployed AssemAI.
@@ VIII Conclusion and Further Work
In this study, we derived an industry-standard dataset tailored for assembly processes using the novel YOLO-FF model and introduced AssemAI, a standardized image-based anomaly detection pipeline. We began with a simple CNN Model, which provided a baseline for anomaly detection through a straightforward architecture and standard training techniques. Building on this, we explored ViT and EfficientNet models to enhance classification accuracy and efficiency. A significant aspect of our approach is the focus on interpretability, where we aimed to understand both model and user lever explainability to high-level phenomena such as structural integrity. This not only improved model understanding but also provided actionable insights for anomaly detection. Our findings indicate that EfficientNet offers significant improvements over traditional methods. Future work should explore hybrid architectures, which take other modalities like time series and textual, further interpretability techniques and deployment on edge devices for real-time anomaly detection applications to extend these findings to other domains and production environments. Also, to improve how domain experts understand our approach, we suggest creating abstract representations of causal factors associated with anomalies. This approach focuses on linking sensor data to broader concepts like structural issues or gripper malfunctions, providing a more comprehensive view of anomalous events.
@@ Acknowledgment
This work is supported in part by NSF grants #2119654, “RII Track 2 FEC: Enabling Factory to Factory (F2F) Networking for Future Manufacturing” and SCRA grant “Enabling Factory to Factory (F2F) Networking for Future Manufacturing across South Carolina”.